OpenStudy (lgbasallote):

solve by linear DE \[(x-y)dx - (x+y)dy =0\]

OpenStudy (lgbasallote):

i think i divide by \[\frac{1}{(x+y)dx}?\]

OpenStudy (lgbasallote):

so i have \[\frac{x-y}{x+y} - \frac{dy}{dx}\]

OpenStudy (lgbasallote):

\[\frac{x}{x+y} - \frac{y}{x+y} = \frac{dy}{dx}\] how do i solve this by linear DE?

OpenStudy (anonymous):

let y=ux and u is a function of x

OpenStudy (lgbasallote):

homogeneous?

OpenStudy (anonymous):

yep thats right

OpenStudy (lgbasallote):

cant it be solved by LDE?

OpenStudy (anonymous):

i think after sub u will have a LDE

OpenStudy (lgbasallote):

ahh i see..thanks for the hint ^_^